We realized that the SRM sensors are connected to the readout but just sitting on the BS in vacuum table with no magnets and therefore no shadows in them. We swapped the inputs to the SRM and PRM satellite boxes to use the higher transimpedance gain of the PRM side sensor. The attached plot shows the current spectrum in this configuration. The PD readback voltage was 9.5 V. Since this is close to the rail we put a slightly higher voltage into the AA of this channel to test that we can read out more ADC counts to make sure we are not saturating. The margin was 15800 vs 15400 counts with p-p of 5 counts on the dataviewer 1 second trend. We returned all cables to nominal configuration.
The calibration from A to m is 59 uA/1 mm.